OpenStudy (anonymous):

any help? Evaluate the integral ∫e^2t sin(-5t) dt ?

OpenStudy (anonymous):

\[\int\limits e^(2t) \sin(-5t) dt\]

OpenStudy (turingtest):

This will be a "repeating" integral. Integrate by parts.

OpenStudy (anonymous):

i tried that it didnt work?

OpenStudy (anonymous):

Use Euler's equation to simplify if you've covered that.

OpenStudy (anonymous):

By a "repeating" integral, what Turing means is that integrate by parts (the integral that you see when you integrate the first time) a second time. When you integrate by parts a second time, you'll notice that the integrals eventually cancel and you get something nice.

OpenStudy (anonymous):

\[\Im(\int\limits e^{(2-5i)t}dt)=\Im(\frac{1}{2-5i}e^{(2-5i)t})\]

OpenStudy (anonymous):

i did it this way: \[e^(2t) \int\limits \sin(-5t) dt \]

OpenStudy (turingtest):

t is not a constant so you cannot take it out of the integral do you know how to integrate by parts?

OpenStudy (anonymous):

but its 2xt?

OpenStudy (anonymous):

Whether it's 2 times t or just t, t is a variable, not a constant. You cannot remove \[e ^{2t}\] from the integral as a constant since you are integrating with respect to t.

OpenStudy (anonymous):

i still cant get it?

OpenStudy (anonymous):

Here is what you are looking for, a step-by-step calculation, using integration by parts two times, successively: http://mathforum.org/library/drmath/view/53749.html

OpenStudy (anonymous):

still confused?

OpenStudy (anonymous):

\[\int e^{2t}\sin(-5t)\;dt\] Since sine is an odd function, you have sin(-5t) = -sin(5t) \[-\int e^{2t}\sin(5t)\;dt\] Integrating by parts, let \[\begin{matrix}u=e^{2t}& & dv=\sin(5t)\;dt\\ du=2e^{2t}dt& & v=-\frac{1}{5}\cos(5t)\end{matrix}\\ \begin{align*}-\int e^{2t}\sin(5t)\;dt&=-\frac{1}{5}e^{2t}\cos(5t)-\int\left(-\frac{1}{5}\cos(5t)\right)\left(2e^{2t}dt\right)\;dt\\ &=-\frac{1}{5}e^{2t}\cos(5t)+\frac{2}{5}\int e^{2t}cos(5t)\;dt\end{align*}\] For the next integral, repeat integration by parts: \[\begin{matrix}f=e^{2t}& & dg=\cos(5t)\;dt\\ df=2e^{2t}dt& & g=\frac{1}{5}\sin(5t)\end{matrix}\\ \begin{align*}-\int e^{2t}\sin(5t)\;dt&=-\frac{1}{5}e^{2t}\cos(5t)+ \\&\;\;\;\;\;\;\;\;\;\; \frac{2}{5}\left[\frac{1}{5}e^{2t}\sin(5t)-\int\left(\frac{1}{5}\sin(5t)\right)\left(2e^{2t}\right)\;dt\right]\\ \color{red}{-\int e^{2t}\sin(5t)\;dt}&=-\frac{1}{5}e^{2t}\cos(5t)+ \\&\;\;\;\;\;\;\;\;\;\; \frac{2}{25}e^{2t}\sin(5t)\color{red}{-\frac{4}{25}\int e^{2t}\sin(5t)\;dt}\end{align*}\] The red text is the repetition @TuringTest was referring to.

OpenStudy (anonymous):

so if i work that out it should be the answer? ?

OpenStudy (anonymous):

Yes. Moving the integral on the RHS to the LHS, then simplifying until you have \[\int e^{2t}\sin(5t)\;dt=\cdots\] Of course, you're supposed to find the antiderivative of e^(2t) sin(-5t), so make sure you rewrite in terms of that.